Lightning occurs when there is a flow of electric charge (principally electrons) between the

ground and a thundercloud. The maximum rate of charge flow in a lightning bolt is about 20,000 C/s this lasts for 100 μs or less. How much charge flows between the ground and the cloud in this time? How many electrons flow during this time?


1
Expert's answer
2022-02-16T10:28:38-0500

"\frac{dq}{dt}=i=20000C/sec"

Charge

"q=it"

"q=20000\times100\times10^{-6}=2C"

Charge

"q=2C"

Number of electrons

"q=ne"

"n=\frac{q}{e}"

"n=\frac{2}{1.6\times10^{-19}}=1.25\times10^{19}"

Number of electrons (n)="1.25\times10^{19}"